OpenStudy (anonymous):

HELP!!! VERY URGENT!!! Find the derivative of f(x) = (x^2 - 4x)(√x - 1).

OpenStudy (anonymous):

\[(x^2 - 4x)(\sqrt{x - 1} )\] is is your question?

OpenStudy (anonymous):

The √x is by itself.

OpenStudy (anonymous):

\[\left( x^{2}-4 \right)\left( \sqrt{x}-1 \right)\]

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

(5/2)x^(3/2) - 2x - 6x^(1/2) + 4

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }x ^{\frac{ 3 }{2 }} - 2x - 6\sqrt{x} + 4\]

OpenStudy (jdoe0001):

$$\bf \left( x^{2}-4 \right)\left( \sqrt{x}-1 \right)\\ \implies \left( x^{2}-4 \right)\left( x^{\frac{1}{2}}-1 \right) $$ then use the product rule

OpenStudy (anonymous):

Could the answer also be , \[\sqrt{x}\left( \frac{ 5}{ 2}x - 6\right)-2x+4\]

OpenStudy (anonymous):